# Twists of Superconformal Algebras

### v2

@@ 3.3.2. Topological Field Theories
In contrast, we can instead consider an affine subspaces $\mathbb{R}^{k}\hookrightarrow C(\mathbb{R}^{p,q})$ on which the $\mathcal{Q}$-twisted theory is fully topological, i.e., the entire algebra of translations of $\mathbb{R}^{k}$ is $\mathcal{Q}$-exact. Examples of this nature generalize the topological twists of supersymmetric field theories first studied by Witten [Witten_1988], but conformal twists allow us to expand the range of possible supercharges to which we may apply the twisting construction.
The structure of the local observables in a topological twist can be described naturally using the structure of an $\mathbb{E}_{k}$-algebra. Let us briefly overview the idea behind these structure, though we will not give a complete definition here (see for instance [CG1, Chapter 3.6.4] and [MarklShniderStasheff, Chapter 4.1] for more details and context).
Let $A$ be a complex vector space, or maybe more generally a cochain complex. To give $A$ the structure of an $\mathbb{E}_{k}$-algebra is to define a multiplication operator $m_{e}\colon A^{\otimes j}\to A$ for each choice of embedding
@@
where $B_{k}(0,1)$ is the $k$-dimensional open unit ball in $\mathbb{R}^{k}$, and $B_{k}(x_{i},r_{i})$ is a sub-ball of radius $r_{i}$ and center $x_{i}$. These multiplication operators should vary smoothly as we vary the $x_{i}$ and $r_{i}$, and they should be “associative,” meaning compatible under concentric inclusions of $k$-balls.
@@
There is an enhanced version of an $\mathbb{E}_{k}$-algebra called a framed $\mathbb{E}_{k}$-algebra: instead of embeddings of $k$-balls, one studies framed embeddings. This amounts to keeping track of the action of the rotation group $\mathrm{SO}(k)$ on $A$. We will not go into detail here, except to point out that a framed $\mathbb{E}_{k}$-algebra may be used to construct a topological field theory on a general oriented $k$-manifold by the process of factorization homology as developed by Ayala and Francis [AyalaFrancis]. See [ScheimbauerThesis] for a full development of this idea.
@@ 3.3.3. Higher Dimensional Non-Topological Examples
Finally, let us briefly consider a more expansive class of examples, where we study an affine patch of dimension $>2$ that is nevertheless not fully topological. Holomorphic examples occur when there is an even real-dimensional affine subspace that we identify as a copy of $\mathbb{C}^{d}\subseteq C(\mathbb{R}^{p,q})$, all of whose isometries are $\mathcal{Q}$-closed and whose anti-holomorphic complexified isometries are $\mathcal{Q}$-exact. One may also study affine subspace of the form $\mathbb{C}^{d}\times\mathbb{R}^{k}\subseteq C(\mathbb{R}^{p,q})$ where the $\mathcal{Q}$-twist is holomorphic in the first factor — in the sense we just outlined — and topological in the second factor — in the sense of the previous subsection.
Let us say more about one specific example, namely the case of an affine patch of the form ×$\mathbb{C}\times\mathbb{R}\hookrightarrow C(\mathbb{R}^{p,q})$ with one holomorphic dimension and one topological dimension. Let us use $z$ for a holomorphic coordinate on $\mathbb{C}$ and $t$ for a coordinate on $\mathbb{R}$. Suppose we find a twisting supercharge $\mathcal{Q}$ satisfying the following conditions:
@@
We expect then — under the function-analytic conditions of Theorem 8.15 in the $\mathbb{C}$ direction — that the local observables in the $\mathcal{Q}$-twisted theory will have the structure of a raviolo vertex algebra in the sense recently described by Garner and Williams [GWraviolo].
@@ 3.4. Connections to the Physics Literature
Twists of supersymmetric theories have been studied for a long time. One of the first examples of twists, giving rise to “topological” theories were introduced with the work of Witten [Witten_1988], which, as we mentioned, ignited the quantum field theoretic interpretation of Donaldson invariants of 4-manifolds. The simplest supersymmetric twists on $\mathbb{R}^{4}$ by a single nilpotent supercharge $\mathcal{Q}$ gives rise to a purely topological theory, since all the translation generators $P^{\mu}$ are in the exact subalgebra $\mathfrak{b}_{\mathcal{Q}}$.
@@
In superconformal theories, there are more complicated examples, identified by classes of square-zero elements $\mathcal{Q}$ in the superconformal algebra. These can be general combinations of ordinary and conformal supercharges. Thus, there is a more complicated structures of $\mathcal{Q}$-closed and $\mathcal{Q}$-exact generators in the even part $\mathfrak{sconf}(n|S_{\mathbb{C}},\mathbb{C})_{0}$, giving rise to more interesting subvarieties where the protected sector of operators lives.
Proposition 3.15.
@@
The type of generators that appear in the closed and exact subalgebras determine how the twisted theory will look like and behave.
@@
A conformal twist of 4d $\mathcal{N}{=}2$ superconformal theories, described in [Beem_2015] has been the launching pad for several works works looking at algebraic structures corresponding to protected sectors of various superconformal theories. In this example, the nilpotent element is a combination of a super Poincaré charge and a superconformal one, so it can only exists in superconformal theories (but not in supersymmetric ones) The structure of the closed and exact subalgebras implies that only some of the translations are closed and/or exact, giving rise to an interesting “mixed” structure, where local operators from the 4d theory are restricted to a 2d complex plane inside space-time. The holomorphic coordinate on this plane corresponds to a closed (complex) translation, whereas the antiholomorphic coordinate, together with one of the real transverse translations, are in the exact subalgebra. The quotient $\mathfrak{z}_{\mathcal{Q}}/\mathfrak{b}_{\mathcal{Q}}$ is an $\mathfrak{sl}_{2}$ algebra, that is then extended to a Virasoro algebra by identifying a candidate stress-energy tensor, whose modes $L_{m}$ give rise to an infinite chiral symmetry algebra. The restriction of local operators to the complex plane and the action of this algebra provides a correspondence between (protected sectors of) local operators of $24d\mathcal{N}{=}$ 2 SCFTs and 2d vertex operator algebras.
Since this seminal work, the standard lore in the community has been that the correspondence between $24d~{}\mathcal{N}{=}$ 2 SCFTs and vertex operator algebras is one-to-one, modulo restricting to only local operators [Argyres:2022npi] or (not) allowing further restrictions of the spectrum [Buican:2023efi].
@@
If one considers extended operators in this setting, it is also important to consider that these operators can stretch to infinity. In this case the signature of the underlying space-time may have a considerable impact on the behavior of the twisted theory, compared to the case when one only considers local operators [Argyres:2022npi]. We will not focus on these subtle features in the present paper and work as much as possible with complex algebras and the largest embedding spaces possible. It is worth noting, however, that when writing explicitly a physical theory, these subtleties ought to be addressed.
Table 1 provides a sample of the conformal twists that have appeared in the physics literature. We exclude those twists of SCFTs that are generated by nilpotent elements constructed purely from Poincaré supercharges, since these are better known and have already been classified (for a similar table of examples in the super Poincaré case, see Tables 1–4 of [ElliottSafronovWilliams]).
@@
3d $\mathcal{N}=4$ on 1d protected associative algebra ×$\mathbb{R}\times\mathbb{R}^{2}$ [Beem:2016cbd]
4d $\mathcal{N}=2$ on $\mathbb{R}^{4}$ 2d Vertex Operator Algebra $\mathbb{C}\sim\mathbb{R}^{2}\subseteq\mathbb{R}^{4}$ [Beem_2015]
4d $\mathcal{N}=2$ on $\mathbb{R}^{1,3}$ 2d Extended VOA $\mathbb{C}\sim\mathbb{R}^{2}\subseteq\mathbb{R}^{1,3}$ [Argyres:2022npi]
6d $\mathcal{N}=(2,0)$ on $\mathbb{R}^{6}$ 2d VOA $\mathbb{C}\sim\mathbb{R}^{2}\subseteq\mathbb{R}^{6}$ [Beem_2015c]
@@
In this section we explore the 4-dimensional case. We begin by describing each complex nilpotence variety and then analyze its orbits under the action of the associated complex superconformal group, including the closed and exact transformations for many twists. These results are available in [DufloSerganova] but we give the arguments here, both to demonstrate how concrete and accessible they are and also to describe them in a way convenient for later results.
@@ 4.1. Complex Nilpotence Variety for $\mathcal{N}\neq 4$
$\cong S_{+}\otimes W_{k}\oplus S_{-}\otimes W_{k}^{*}$
where $W_{k}$ denotes the defining $k$-dimensional representation of the factor $\mathfrak{sl}(k,\mathbb{C})$ and $V_{4}$ denotes the four-dimensional defining representation of $\mathfrak{sl}(4,\mathbb{C})$, which is isomorphic to the positive semispin (or Weyl spinor) representation $S_{+}$ of $\mathfrak{so}(6,\mathbb{C})$. The one-dimensional factor in the center of $(\mathfrak{sconf}(4|k,\mathbb{C}))_{0}$ acts trivially.
@@
Any odd element decomposes into a sum of terms from each summand. A pure tensor in the first summand has the form $Q_{+}\otimes w_{+}$ with $Q_{+}\in S_{+}$ and $w_{+}\in W_{k}$. A pure tensor in the second summand has the form $Q_{-}\otimes w_{-}$ with $Q_{+}\in S_{+}$ and $w_{-}\in W_{k}^{*}$. Pairs of elements in the same summand of $(\mathfrak{sconf}(4|k,\mathbb{C}))_{1}$ have trivial bracket. The Lie bracket between a pair of elements in opposite summands $[Q_{+}\otimes w_{+},Q^{\prime}_{-}\otimes w^{\prime}_{-}]$ is given by
@@
$[\mathcal{Q},\mathcal{Q}]=2[\mathcal{Q}_{+},\mathcal{Q}_{-}],$
@@ 4.2. The $\mathcal{N}=4$ Case
$(\mathfrak{sconf}(4|k,\mathbb{C}))_{1}\cong V_{4}\otimes W_{4}\oplus V_{4}^{*}\otimes W_{4}^{*}\cong S_{+}\otimes W_{4}\oplus S_{-}\otimes W_{4}^{*}$
@@
Let us discuss the subspace of this nilpotence variety consisting of nilpotent elements in the super Poincaré algebra associated to a choice of embedding $\mathfrak{so}(4,\mathbb{C})\subseteq\mathfrak{so}(6,\mathbb{C})$. Under this embedding the representation $V_{4}$ decomposes as $S_{+}\oplus S_{-}$, where $S_{\pm}$ are the two two-complex-dimensional semispin representations of $\mathfrak{so}(4,\mathbb{C})$. The complex nilpotence variety for the $\mathcal{N}=4$ super Poincaré algebra is isomorphic to
@@
where $\mathcal{Q}^{\prime}_{+}\in S_{+}\subseteq V_{4}$ and $\mathcal{Q}^{\prime}_{-}\in S_{-}\cong S_{-}^{*}\subseteq V_{4}^{*}$. Note that the condition $\mathrm{Tr}(Q^{\prime}_{+}\otimes Q^{\prime}_{-})=0$ is automatically satisfied. The decomposition of this super Poincaré nilpotence variety into $\mathrm{Spin}(4,\mathbb{C})\times\mathrm{SL}(4,\mathbb{C})$-orbits is interesting, including for instance the one-parameter families of topological supercharges studied by Kapustin and Witten [KapustinWitten]. This decomposition is discussed in [ElliottSafronov, Section 4.4] and [EagerSaberiWalcher, Section 9].

### v3

@@ 3.3.2. Topological Field Theories
In contrast, we can instead consider an affine subspaces $\mathbb{R}^{k}\hookrightarrow C(\mathbb{R}^{p,q})$ on which the $\mathcal{Q}$-twisted theory is fully topological, i.e., the entire algebra of translations of $\mathbb{R}^{k}$ is $\mathcal{Q}$-exact. Examples of this nature generalize the topological twists of supersymmetric field theories first studied by Witten [45], but conformal twists allow us to expand the range of possible supercharges to which we may apply the twisting construction.
The structure of the local observables in a topological twist can be described naturally using the structure of an $\mathbb{E}_{k}$-algebra. Let us briefly overview the idea behind these structure, though we will not give a complete definition here (see for instance [15, Chapter 3.6.4] and [35, Chapter 4.1] for more details and context).
Let $A$ be a complex vector space, or maybe more generally a cochain complex. To give $A$ the structure of an $\mathbb{E}_{k}$-algebra is to define a multiplication operator $m_{e}\colon A^{\otimes j}\to A$ for each choice of embedding of $j$ disjoint $k$-balls into a bigger $k$-ball:
@@
with $B_{k}(0,1)$ the $k$-dimensional open unit ball in $\mathbb{R}^{k}$, and $B_{k}(x_{i},r_{i})$ is a sub-ball of radius $r_{i}$ and center $x_{i}$. These multiplication operators should vary smoothly as we vary the centers $x_{i}$ and radii $r_{i}$, and they should be “associative,” meaning compatible under (rescaled) inclusions of $k$-balls.
@@
There is a variant version of an $\mathbb{E}_{k}$-algebra called a framed $\mathbb{E}_{k}$-algebra relevant to oriented manifolds: instead of embeddings of $k$-balls, one studies framed embeddings. This amounts to keeping track of the action of the rotation group $\mathrm{SO}(k)$ on $A$. We will not go into detail here, except to point out that a framed $\mathbb{E}_{k}$-algebra may be used to construct a topological field theory on a general oriented $k$-manifold by the process of factorization homology as developed by Ayala and Francis [4]. See [41] for a full development of this idea.
@@ 3.3.3. Higher Dimensional Non-Topological Examples
Finally, let us briefly consider a more expansive class of examples, where we study an affine patch of dimension >2>2 that is nevertheless not fully topological. Holomorphic examples occur when there is an even real-dimensional affine subspace that we identify as a copy of $\mathbb{C}^{d}\subseteq C(\mathbb{R}^{p,q})$, all of whose isometries are $\mathcal{Q}$-closed and whose anti-holomorphic complexified isometries are $\mathcal{Q}$-exact. One may also study affine subspace of the form $\mathbb{C}^{d}\times\mathbb{R}^{k}\subseteq C(\mathbb{R}^{p,q})$ where the $\mathcal{Q}$-twist is holomorphic in the first factor — in the sense we just outlined — and topological in the second factor — in the sense of the previous subsection.
Let us say more about one specific example, namely the case of an affine patch of the form $\mathbb{C}\times\mathbb{R}\hookrightarrow C(\mathbb{R}^{p,q})$ with one holomorphic dimension and one topological dimension. Let us use $z$ for a holomorphic coordinate on $\mathbb{C}$ and $t$ for a coordinate on $\mathbb{R}$. Suppose we find a twisting supercharge $\mathcal{Q}$ satisfying the following conditions:
@@
We expect then — under the function-analytic conditions of Theorem 8.15 in the $\mathbb{C}$ direction — that the local observables in the $\mathcal{Q}$-twisted theory will have the structure of a raviolo vertex algebra in the sense recently described by Garner and Williams [26]. See [39] for a discussion in the language of Poisson vertex algebras, with many physical applications.
3.3.4. Reality Conditions
The superconformal Lie groups and algebras have real forms as well as complexifications, and it is useful to track what phenomena depend upon (or manifest) the real form or the complex form.
For a given real twisting supercharge $\mathcal{Q}$, we can study its real Lie algebras of closed and exact elements $\mathfrak{z}_{\mathcal{Q}}$ and $\mathfrak{b}_{\mathcal{Q}}$. These have geometric significance: these real vector fields generate flows, and $\mathcal{Q}$-closed observables are locally constant along these flows. This feature leads to strong constraints on expected values, correlation functions, and operator product expansions. As a rich but rather simple example, we explore 4d $\mathcal{N}=2$ chiral supercharges in Section 5.3.1. It would be interesting to explore more examples, but we do not address any in the present work.
More generally, for any twisting supercharge, there is still interesting structure on the observables of a twisted theory coming from closed and exact complex Lie algebra elements. As the observables form a complex vector space, it is natural to work with this complexification of the superconformal Lie algebra. The $\mathcal{Q}$-closed observables are now invariant under the action of $\mathcal{Q}$-closed or exact elements of the superconformal algebra, and such a condition can still lead to nontrivial and powerful constraints. We saw above in Section 3.3.1, for example, how $\partial/\partial\bar{z}$ can appear as a $\mathcal{Q}$-exact element and how that imposes holomorphicity on the $\mathcal{Q}$-twisted observables, leading to a vertex algebra. This kind of phenomenon — an analogue of holomorphicity — is available in many of the situations we explore in this paper but to fully characterize it requires developing analogues of vertex algebras in novel situations, a problem well worth studying.
@@ 3.4. Connections to the Physics Literature
Twists of supersymmetric theories have been studied for a long time. One of the first examples of twists, giving rise to a “topological” theory, was introduced with the work of Witten [45], which provided a quantum field theoretic interpretation of Donaldson invariants of 4-manifolds. This twist on $\mathbb{R}^{4}$ is a nilpotent supercharge $\mathcal{Q}$ where all the translation generators $P^{\mu}$ are in the exact subalgebra $\mathfrak{b}_{\mathcal{Q}}$ and hence gives rise to a purely topological theory.
@@
In superconformal theories, there are more complicated examples of square-zero odd elements $\mathcal{Q}$ in the superconformal algebra, involving linear combinations of supersymmetric and conformal supercharges. Thus, there is a more complicated structures of $\mathcal{Q}$-closed and $\mathcal{Q}$-exact generators in the even part $\mathfrak{sconf}(n|S,\mathbb{C})_{0}$, giving rise to more interesting subvarieties where the protected sector of operators lives. For a physicist, the essential point is this:
@@
The type of generators that appear in the closed and exact subalgebras determine how the twisted theory will look and behave.
@@
As an example, there is a conformal twist of 4d $\mathcal{N}{=}2$ superconformal theories, described in [6], that has been the launching pad for a slew of works studying algebraic structures corresponding to protected sectors within superconformal theories. In this example, the nilpotent element is a combination of a super Poincaré charge and a superconformal charge, so it only exists in superconformal theories (but not in supersymmetric ones). The structure of the closed and exact subalgebras implies that only some of the translations are closed and only some are exact, giving rise to an interesting “mixed” structure, where local operators from the 4d theory are restricted to a 2d complex plane in space-time. The holomorphic coordinate on this plane corresponds to a closed (complex) translation, whereas the antiholomorphic coordinate, together with one of the real transverse translations, are in the exact subalgebra. The quotient $\mathfrak{z}_{\mathcal{Q}}/\mathfrak{b}_{\mathcal{Q}}$ is an $\mathfrak{sl}_{2}$ algebra, that is then extended to a Virasoro algebra by identifying a candidate stress-energy tensor, whose modes $\{L_{m}\}$ give rise to an infinite chiral symmetry algebra. The restriction of local operators to the complex plane and the action of this algebra provides a correspondence between (protected sectors of) local operators of $4d\,\mathcal{N}{=}2$ SCFTs and 2d vertex operator algebras.
Since this seminal work, the lore has been that the correspondence between $4d~\mathcal{N}{=}2$ SCFTs and vertex operator algebras is one-to-one: that is, given a vertex operator algebra, one can reconstruct the parent SCFT. This claim, however, cannot be generically true. The $4d~\mathcal{N}{=}4$ theories have a conformal manifold, and each point in it corresponds to a different SCFT, while the VOA is insensitive to exactly marginal couplings. In fact, the VOA obtained after the twist is the same for all points in the conformal manifold. Furthermore, the VOA data is only sensitive to the local operator data. Again in the case of $4d~\mathcal{N}{=}4$, theories are distinguished by global forms of gauge groups, which depend on the spectrum of line operators [2]. (These perturbative descriptions live at corners of the aforementioned conformal manifolds.) Furthermore, by including line operators in the conformal twist of $4d~\mathcal{N}{=}2$ SCFTs (new cohomology classes relative to the ones obtained from local operators), one obtains a bigger, richer vertex algebra structure that contains the local operator VOA [3]. The map from 4d SCFT to VOA can be further modified by imposing additional restrictions on the local operators allowed after the twist, and this gives rise to a vertex algebra that is a “square root” of the original VOA [13].
@@
If one considers extended operators in this setting, it is also important to consider that these operators can stretch to infinity. In this case the signature of the underlying space-time may have a considerable impact on the behavior of the twisted theory, compared to the case when one only considers local operators [3]. We will not focus on these subtle features in the present paper and work as much as possible with complex algebras and the largest embedding spaces possible. It is worth noting, however, that when writing explicitly a physical theory, these subtleties ought to be addressed.
Enlarging the amount of supersymmetry, relative to the Beem et al. case, gives rise to additional examples, notably the Kapustin-Witten twist[32]. In this case, although the nilpotent element is a combination of Poincaré supercharges, an identification with the enlarged R-symmetry is used, giving rise to a 2d topological theory. Other types of twists of $\mathcal{N}{\geq}2$ SCFTs have been introduced by putting the theory on a non-trivial background and twisting by it (i.e., an $\Omega$-background) effectively localizing the theory on a lower dimensional locus in space-time [38]. Nevertheless, in the case of 4d $\mathcal{N}{=}4$ SYM, a twist of this kind can be equivalently realized without specifying any background, by simply using a particular combination of supercharges. This localizes the theory on a 2d slice, as in the Beem et al. case, but with an additional 2d transverse quasi-topological sector [17].
Table 1 provides a sample of the conformal twists that have appeared in the physics literature. See [17] for a more thorough discussion (and many pertinent references). We exclude those twists of SCFTs that are generated by nilpotent elements constructed purely from Poincaré supercharges, since these are better known and have already been classified (for a similar table of examples in the super Poincaré case, see Tables 1–4 of [22]).
@@
3d $\mathcal{N}=4$ on $\mathbb{R}^{3}$ 1d protected associative algebra $\mathbb{R}\times\mathbb{R}^{2}$ [8]
4d $\mathcal{N}=2$ on $\mathbb{R}^{4}$ 2d Vertex Operator Algebra $\mathbb{C}\sim\mathbb{R}^{2}\subseteq\mathbb{R}^{4}$ [6]
4d $\mathcal{N}=4$ on $\mathbb{R}^{4}$ A/B 2d TQFTs $\mathbb{C}\times\Sigma$ [32]
4d $\mathcal{N}=4$ on $\mathbb{R}^{4}$ A/B 2d quasi-topological sector $\mathbb{C}\times\mathbb{R}^{2}$ [17]
4d $\mathcal{N}=2$ on $\mathbb{R}^{1,3}$ 2d Extended VOA $\mathbb{C}\sim\mathbb{R}^{2}\subseteq\mathbb{R}^{1,3}$ [3]
6d $\mathcal{N}=(2,0)$ on $\mathbb{R}^{6}$ 2d VOA $\mathbb{C}\sim\mathbb{R}^{2}\subseteq\mathbb{R}^{6}$ [9]
@@
In this section we explore the 4-dimensional case. We begin by describing each complex nilpotence variety and then analyze its orbits under the action of the associated complex superconformal group, including the closed and exact transformations for many twists. These results are available in [20] but we give the arguments here, both to demonstrate how concrete and accessible they are and also to describe them in a way convenient for later results.
@@ 4.1. Complex Nilpotence Variety for $\mathcal{N}\neq 4$
$\cong S^{(6)}_{+}\otimes W_{k}\oplus S^{(6)}_{-}\otimes W_{k}^{*}$
where $W_{k}$ denotes the defining $k$-dimensional representation of the factor $\mathfrak{sl}(k,\mathbb{C})$ and $V_{4}$ denotes the four-dimensional defining representation of $\mathfrak{sl}(4,\mathbb{C})$, which is isomorphic to the positive semispin (or Weyl spinor) representation $S^{(6)}_{+}$ of $\mathfrak{so}(6,\mathbb{C})$. The one-dimensional factor in the center of $(\mathfrak{sconf}(4|k,\mathbb{C}))_{0}$ acts trivially on $(\mathfrak{sconf}(4|k,\mathbb{C}))_{0}$ but it acts nontrivially on $(\mathfrak{sconf}(4|k,\mathbb{C}))_{1}$, as we will see below.
@@
Any odd element decomposes into a sum of terms from each summand. A pure tensor in the first summand has the form $Q_{+}\otimes w_{+}$ with $Q_{+}\in S^{(6)}_{+}$ and $w_{+}\in W_{k}$. A pure tensor in the second summand has the form $Q_{-}\otimes w_{-}$ with $Q_{+}\in S^{(6)}_{+}$ and $w_{-}\in W_{k}^{*}$. Pairs of elements in the same summand of $(\mathfrak{sconf}(4|k,\mathbb{C}))_{1}$ have trivial bracket. The Lie bracket between a pair of elements in opposite summands $Q_{+}\otimes w_{+},Q^{\prime}_{-}\otimes w^{\prime}_{-}]$ is given by
@@
$\mathcal{Q},\mathcal{Q}]=2[\mathcal{Q}_{+},\mathcal{Q}_{-}],$
@@ 4.2. The $\mathcal{N}=4$ Case
$(\mathfrak{sconf}(4|k,\mathbb{C}))_{1}\cong V_{4}\otimes W_{4}\oplus V_{4}^{*}\otimes W_{4}^{*}\cong S^{(6)}_{+}\otimes W_{4}\oplus S^{(6)}_{-}\otimes W_{4}^{*}$
@@
Let us discuss the subspace of this nilpotence variety consisting of nilpotent elements in the super Poincaré algebra associated to a choice of embedding $\mathfrak{so}(4,\mathbb{C})\subseteq\mathfrak{so}(6,\mathbb{C})$. Under this embedding the representation $V_{4}$ decomposes as $S^{(4)}_{+}\oplus S^{(4)}_{-}$, where $S^{(4)}_{\pm}$ are the two semispin representations of $\mathfrak{so}(4,\mathbb{C})$ (each of complex dimension 2). The complex nilpotence variety for the $\mathcal{N}=4$ super Poincaré algebra is isomorphic to
@@
where $\mathcal{Q}^{\prime}_{+}\in S^{(4)}_{+}\subseteq S^{(6)}_{+}\cong V_{4}$ and $\mathcal{Q}^{\prime}_{-}\in S^{(4)}_{-}\subseteq S^{(6)}_{-}\cong V_{4}^{*}$. Note that the condition $\mathrm{Tr}(Q^{\prime}_{+}\otimes Q^{\prime}_{-})=0$ is automatically satisfied. The decomposition of this super Poincaré nilpotence variety into $\mathrm{Spin}(4,\mathbb{C})\times\mathrm{SL}(4,\mathbb{C})$-orbits is interesting, including for instance the one-parameter families of topological supercharges studied by Kapustin and Witten [32]. This decomposition is discussed in [23, Section 4.4] and [21, Section 9].